Two particles are seperates by a horizontal distance X and are launched with velocity U√3 at an angle of 30° and another with velocity U at an angle of 60° find the time after which the horizontal distance between them is  0
 

Grade:9

1 Answers

Eshan
5 years ago
Horizontal component of velocity of first particle=\sqrt{3}ucos30^{\circ}=\dfrac{3u}{2}
Horizontal component of velocity of first particle=ucos60^{\circ}=\dfrac{u}{2}


Therefore the velocity of the two particles relative to one another is\dfrac{3u}{2}-\dfrac{u}{2}=u

Hence the time in which the horizontal distance between them is zero is\dfrac{X}{u}